OpenStudy (anonymous):

(cosx)(tanx+sinx cotx)=sinx+cos^2x Prove the identity

OpenStudy (anonymous):

Are you sure it's not (cosx)(tanx+sinx cosx) = sinx + sinx cos²x ?

OpenStudy (anonymous):

oh crap, its actually (cosx)(tanx+sinx cotx)=sinx+cos^2x

OpenStudy (anonymous):

We know that \(\tan x = \dfrac{\sin x}{\cos x}\) and \(\cot x = \dfrac{\cos x}{\sin x}\) So \((\cos x)(\tan x+\sin x \cot x) = (\cos x)\left(\dfrac{\sin x}{\cos x} + \sin x\cdot\dfrac{\cos x}{\sin x}\right)\) \(=(\cos x)\left(\dfrac{\sin x}{\cos x} + \cos x\right)\) \( = \boxed{\sin x + \cos^2 x}\)

OpenStudy (anonymous):

\[\Huge\text{Q.E.D.}\]

OpenStudy (anonymous):

Q.E.D. ? Thanks a lot for the help otherwise

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Q.E.D.

OpenStudy (anonymous):

wait, how'd you get sinx + cos^2 at the end? shouldn't sinx/cosx cross out the + cosx?

OpenStudy (anonymous):

No, because it's just addition. you can't cancel cos x out like that.